OpenStudy (anonymous):

The letters x and y represent rectangular coordinates. Write the following equation using polar coordinates (r, θ). x^2 + y^2 - 4x = 0

OpenStudy (anonymous):

is it rsin^2 tangent = 4cos tangent ?

OpenStudy (anonymous):

@theEric

OpenStudy (theeric):

I don't think so. I won't say I'm 100% sure, because I make mistakes. But let's work through it together. Do you know the first step in converting these?

OpenStudy (anonymous):

welll x= rcos θ y = rsin θ

OpenStudy (theeric):

Right! That's a big part. And so what we do then is substitution. So, check my math, but I'll do just the substitution and get... \(\left(r\cos\theta\right)^2+\left(r\sin\theta\right)^2-4\left(r\cos\theta\right)=0\) How does that look?

OpenStudy (anonymous):

yes thats right

OpenStudy (anonymous):

I got like one step ahead of you but idk if i did it rihgt :/

OpenStudy (theeric):

Okay! Well I'm going to distribute that power as my next step. If we got different equations, we should look at them! I'll type mine out now.

OpenStudy (theeric):

I also took away some parenthesis, where I put it in blue. \(r^2\cos^2\theta+r^2\sin^2\theta-4\color{blue}{r\cos\theta}=0\) So, how do those changes look to you?

OpenStudy (anonymous):

yeah thats what i got too :))

OpenStudy (theeric):

Cool!

OpenStudy (anonymous):

And that is as far as i got as well haha

OpenStudy (theeric):

Okay! You asked "is it rsin^2 tangent = 4cos tangent ?" earlier. You didn't try a next step?

OpenStudy (anonymous):

Thats one of the answers and it looked the closest so i just thought it might be that one haha

OpenStudy (theeric):

That's better than guessing randomly, I suppose! So, the next step! Of all the annoying (difficult to remember) trigonometric equations, there are two I frequently find useful 1. \(\quad\tan\theta=\dfrac{\sin\theta}{\cos\theta}\) and 2. \(\quad\cos^2\theta+\sin^2\theta=1\) We'll use that second one. Usually the \(\cos^2\) and \(\sin^2\) share a common facter that you can distribute out. Is that the case here?

OpenStudy (anonymous):

Kind of our equation equals 0 tho?

OpenStudy (theeric):

The expressions in our equation equal \(0\), yep! "x^2 + y^2 - 4x = 0" and \(r^2\cos^2\theta+r^2\sin^2\theta-4r\cos\theta=0\)

OpenStudy (anonymous):

oh okay so whats the next step??

OpenStudy (theeric):

But we want to simplify what we have. Look at this hint: \(\color{red}{r^2}\color{blue}{\cos^2\theta}+\color{red}{r^2}\color{blue}{\sin^2\theta}-4r\cos\theta=0\) When we see \(\cos^2\theta+\sin^2\theta\) we always want to see if we can isolate it. That's because it will turn into \(1\), which is a lot easier on the eyes! :)

OpenStudy (anonymous):

Does this r^2cos^2θ+r^2sin^2 cancel out because of cos^2θ+sin^2θ=1?

OpenStudy (anonymous):

OK so it become 1 and then what happens to the r^2?

OpenStudy (theeric):

Well, it can't just disappear on its own. You rearrange the equation so that \(\cos^2\theta+\sin^2\theta\) is by itself, and then you'll see how the magic works. So, \(\color{red}{r^2}\color{blue}{\cos^2\theta}+\color{red}{r^2}\color{blue}{\sin^2\theta}-4r\cos\theta=0\) \(\color{red}{r^2}\left(\color{blue}{\cos^2\theta}+\color{blue}{\sin^2\theta}\right)-4r\cos\theta=0\) Right?

OpenStudy (theeric):

Then can you see what happens next?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

r^2- 4rcosθ=0

OpenStudy (theeric):

Right! And then it's up to you whether you want to pull an \(r\) out to change it more or not! Both forms are useful for different things, so I'd say you're done!

OpenStudy (theeric):

I guess just look at the choices and see if it matches now. If not, simplify it one step further, and then check.

OpenStudy (anonymous):

r = 4cos θ is an answer is that one right??

OpenStudy (theeric):

What do you think?

OpenStudy (anonymous):

yes

OpenStudy (theeric):

Yep! Looks good! You add \(4r\cos\theta\) to both sides, and then divide by \(r\). You can divide by \(r\) as long as \(r\neq0\). And if \(r=0\), then you can see that the equation is useless anyway.

OpenStudy (anonymous):

Awesomeee!! :) thanks :D

OpenStudy (theeric):

You're welcome! Good job! :)

OpenStudy (anonymous):

Thx :)

OpenStudy (theeric):

Np!